OpenStudy (anonymous):

Work out the value of this expressions, when p=2, q=-3, r=5: 20-pq^2r/p-q+r

OpenStudy (mimi_x3):

you stil need help, with it ?

OpenStudy (anonymous):

its is another question

OpenStudy (mimi_x3):

its similar to substitute the pronumerals

OpenStudy (mimi_x3):

just substitute*

OpenStudy (anonymous):

does order of operations apply?

OpenStudy (anonymous):

p=2, q=-3, r=5: 20-pq^2r/p-q+r 20-18*5/10 = -70/10 = -7

OpenStudy (anonymous):

meaning working out the division problem then subtract 20

OpenStudy (anonymous):

20-2(9)(5)/2+3+5=7

OpenStudy (anonymous):

it's q square so 18*5=90

OpenStudy (anonymous):

20-90=-70 and -70/10 = -7

OpenStudy (anonymous):

im sorrry i forgot the sign

OpenStudy (anonymous):

there u go hollywood, all done.

OpenStudy (anonymous):

the negative sign

OpenStudy (anonymous):

yes yes, thank you

OpenStudy (anonymous):

np

OpenStudy (anonymous):

actually

OpenStudy (anonymous):

where can I find more of these problems so I can practice,

OpenStudy (anonymous):

where did u get these???

OpenStudy (anonymous):

from my GSCE textbook

OpenStudy (anonymous):

they dont have any more of these?

OpenStudy (anonymous):

no